# Symmetry-enforced Moiré Topology

### v1

@@ V Conclusion and Discussion
The generation of the moiré potential is not limited to twisting two homobilayers—it can also be done (i) by forming hetero-structure with insulating materials such as boron nitride or transition metal dichalcogenides (TMD) [126, 127, 128, 129] and (ii) by fabricating patterned hole arrays in dielectric substrate materials [130, 131]. We note that although the homobilayer structure contains two copies of the $\Gamma$-valley modes from each layer, the low-energy physics can only involve either the bonding combination or the anti-bonding combination between these two copies, owing to the strong uniform interlayer coupling [132, 21, 40, 133]. As a result, at low energies, homobilayer twist structure effectively has only one set of the $\Gamma$-valley modes. This is exactly the case in twisted homobilayer TMD at $\Gamma$ valley [132, 21, 40]. While our current study focuses on layered materials with low-energy $\Gamma$-valley modes, the methodology should be generalizable to other valleys, such as $K$ or $M$ valleys [134], which we leave for the future work.
VI Acknowledgment

### v2

@@ VI Conclusion and Discussion
We note that the TR-breaking cases with the spinless models in Sec.C2 of SM [SM] can also be applied to topological moiré magnon systems [li2020moire, wang2020stacking, ganguli2023visualization]. Furthermore, the generation of the moiré potential is not limited to twisting two homobilayers—it can also be done (i) by forming hetero-structure with insulating materials such as boron nitride or transition metal dichalcogenides (TMD) [yasuda2021stacking, woods2021charge, zhao2021universal, kim2024electrostatic] and (ii) by fabricating patterned hole arrays in dielectric substrate materials [forsythe2018band, barcons2022engineering]. We note that although the homobilayer structure contains two copies of the $\Gamma$-valley modes from each layer, the low-energy physics can only involve either the bonding combination or the anti-bonding combination between these two copies, owing to the strong uniform interlayer coupling [angeli2021gamma, zhang2021electronic, jia2024moire, liu2025ideal]. As a result, at low energies, homobilayer twist structure effectively has only one set of the $\Gamma$-valley modes. This is exactly the case in twisted homobilayer TMD at $\Gamma$ valley [angeli2021gamma, zhang2021electronic, jia2024moire]. While our current study focuses on layered materials with low-energy $\Gamma$-valley modes, the methodology should be generalizable to other valleys, such as $K$ or $M$ valleys [cualuguaru2025moire], which we leave for the future work.
VII Acknowledgment